a triangle has side lengths 1.5cm, 2cm, and 2.5cm. Is the triangle a right, acute, or obtuse triangle?
Answer by checkley77(12569)

1.5^2+2^2=2.5^2 the test for a right triangle.
2.25+4=6.26
6.25=6.25 looks like a right triangle. (a common 3,4,5 right triangle.